OpenStudy (anonymous):

Given u = (-2, -6) and V = 3, -3, find U times V. Do I just multiply them together??

OpenStudy (hba):

So V=(3,-3) ? or V=3,-3

OpenStudy (anonymous):

the first one

OpenStudy (hba):

So there is a rule to multiply ordered pairs : \[ (a,b) * (c,d) = (ac-bd,ad+bc) \]

OpenStudy (hba):

So the Answer is that you have to multiply them.

OpenStudy (phi):

This could be a "dot product" ???

OpenStudy (anonymous):

i got -24 and -2 but idk where to go from here

OpenStudy (hba):

Because if you got that then that must be the answer : (-24,-2) Lemme Check this quick :)

OpenStudy (anonymous):

its not one of my choices though :( i have -24, -6, 12, =12, 24...i got -24 but what would i do with the -2

OpenStudy (phi):

Is there more background on this problem? I would guess it is a dot product \[(a,b) \cdot (c,d)= ac+bd \]

OpenStudy (hba):

@phi Thanks for the feedback that can be it.

OpenStudy (anonymous):

that's probably it, the question just doesn't say much so i was confused on which formula i had to use. let me try it

OpenStudy (phi):

I am sure they have mentioned this in class...

OpenStudy (hba):

@phi What was my formula related to ?

OpenStudy (anonymous):

i got 6..

OpenStudy (phi):

try again

OpenStudy (phi):

@hba if you treat the pairs as real, imaginary you get your formula

OpenStudy (anonymous):

never mind i got 12! i see what i did wrong

OpenStudy (hba):

Thank you. @phi

OpenStudy (phi):

Also, class should have mentioned that \[ u \cdot v= |u| |v| \cos(\theta)\]

OpenStudy (hba):

Vectors :)

OpenStudy (anonymous):

yes I have this formula, but all the notes and example problems that my teacher has given me does not have any problems that are worded with "determine the quantity to indicate the following". The ones he has given me instructs which formula to use..thats why I was confused. Thank you for your help.

OpenStudy (hba):

your'e welcome.